OpenStudy (nessapetrenko):

Solve the equation for y. 5x – 7y = 10

OpenStudy (gabbyalicorn):

Let's solve for x. 5x−7y=10 Step 1: Add 7y to both sides. 5x−7y+7y=10+7y 5x=7y+10 Step 2: Divide both sides by 5. 5x5=7y+105 x=75y+2 Answer: x=75y+2

OpenStudy (nessapetrenko):

it needs to be in a fraction. @gabbyalicorn

OpenStudy (gabbyalicorn):

7/5 

OpenStudy (anonymous):

The first response is incorrect

OpenStudy (anonymous):

Solve for y,,Step one isolate the y variable so we need to get 5x to the other side of the equal sign. what would you do?